OpenStudy (anonymous):

three points a b c are on straight line such that ab=ac. a particled moving with uniform acceleration passes a and c with speeds 6m/s and 8m/s respectively. another particle moving with uniform acceleration passes a and b with speeds of 12m/s and 14m/s respectively it t1 and t2 are the times taken by the particle to cross ab t1/t2 is nearly??

OpenStudy (inkyvoyd):

omg

OpenStudy (anonymous):

what is omg??

OpenStudy (inkyvoyd):

oh-my-god lol

OpenStudy (anonymous):

why??

OpenStudy (inkyvoyd):

Alright. Begin with the equation. \(d=V_0t+(1/2)at^2\)

OpenStudy (anonymous):

ok

OpenStudy (inkyvoyd):

\(\frac{2}{a_1}=t_1\) \(\frac{2}{a_2}=t_2\)

OpenStudy (inkyvoyd):

\(\huge \frac{t_1}{t_2}=\frac{a_2}{a_1}\)

OpenStudy (anonymous):

did nt understand

OpenStudy (anonymous):

2a1=t1 2a2=t2

OpenStudy (inkyvoyd):

(8m/s-6m/s)/2=a1

OpenStudy (inkyvoyd):

same for a2

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

hey from where at 2 came....

OpenStudy (inkyvoyd):

(14m/s-12m/s)/2=a2

OpenStudy (inkyvoyd):

a1 is the accleration of first, a2 is acceleration of second.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

then....

OpenStudy (inkyvoyd):

Now, go back to the first equation.

OpenStudy (anonymous):

hey wait how this came (14m/s-12m/s)/2=a2 a =v2-v1/time

OpenStudy (anonymous):

time is not given

OpenStudy (anonymous):

why did u take it as 2???????

OpenStudy (inkyvoyd):

Sorry, I forgot time.

OpenStudy (inkyvoyd):

my equations in latex should be correct.

OpenStudy (inkyvoyd):

Lol, I'm not completely sure how to do the problem. What's the answer?

OpenStudy (anonymous):

the answer is 2

OpenStudy (anonymous):

three points a b c are on straight line such that ab=ac. a particled moving with uniform acceleration passes a and c with speeds 6m/s and 8m/s respectively. another particle moving with uniform acceleration passes a and b with speeds of 12m/s and 14m/s respectively it t1 and t2 are the times taken by the particle to cross ab t1/t2 is nearly??

OpenStudy (inkyvoyd):

The key is "is nearly". we have to approximate the values, and I'm not sure how to do that.

OpenStudy (anonymous):

leave it next question.....